### v1

@@ 2.2 Metrics for LLM Reasoning Capacity Boundary
For coding tasks, where a compiler and predefined unit test cases are used as verifiers, the pass@$k$ value can accurately reflect whether the model can solve the problem. However, the issue of “hacking” can become pronounced in mathematics as $k$ increases. In math problems, a model may generate an incorrect chain of thought but still accidentally arrive at the correct answer in one of many samplings, a scenario that is often overlooked by previous metrics. To address this, we filter out problems that are easily “hacked” and then manually check the correctness of CoT for a subset of model outputs. By combining these measures, we rigorously evaluate the limit of LLM’s reasoning capacity as detailed in Section 3.1.
Table 1: Experimental setup for assessing RLVR’s effect on the reasoning boundaries of LLMs across different tasks.
Task Start Model RL Framework RL Algorithm(s) Benchmark(s)
Mathematics LLaMA-3.1-8B Qwen-2.5-7B/14B/32B-Base Qwen-2.5-Math-7B SimpleRLZoo Oat-Zero GRPO GSM8K, MATH500 Minerva, Olympiad AIME24, AMC23
Code Generation Qwen-2.5-7B-Instruct Code-R1 GRPO      LiveCodeBench HumanEval+
Visual Reasoning Qwen-2.5-VL-7B EasyR1 GRPO MathVista MathVision
Deep Analysis Qwen-2.5-7B-Base Qwen-2.5-7B-Instruct DeepSeek-R1-Distill-Qwen-7B VeRL PPO, GRPO Reinforce++ RLOO, ReMax, DAPO     Omni-Math-Rule MATH500
3 RLVR’s Effect on Reasoning Capacity Boundary
Evaluation Protocal. For all sampling procedures involving both base and RL-trained models, we use a temperature of 0.6 and a top-$p$ value of 0.95, allowing a maximum generation of 16,384 tokens. We also show the effect of different temperature settings in Figure 10. For evaluation of the base model, a common practice is to include few-shot examples in the prompt to guide both the output formatting and the reasoning process Grattafiori et al. (2024); Yang et al. (2024); Liu et al. (2024). However, to ensure a fair and unbiased comparison, we deliberately avoid using few-shot prompts for base models, eliminating any potential confounding effects on reasoning that might be introduced by in-context examples. For evaluating both the base and RLVR models, we use the same zero-shot prompt as in RLVR training, or the default prompt provided by the benchmark, ensuring a consistent setup across both models. Interestingly, although base models often produce unformatted or non-sensical responses without few-shot guidance, we observe that with sufficient sampling, they are still capable of generating correctly formatted outputs and successfully solving complex problems. Prompt templates for training and evaluation are provided in Appendix B.
3.1 RLVR for Mathematical Reasoning
Models and Benchmarks. In math problems, models are required to generate a reasoning process (i.e., CoT) along with the final answer. To ensure the robustness of our conclusions, we experiment with multiple LLM families, primarily Qwen-2.5 (7B/14B/32B base variants) Yang et al. (2024) and additionally LLaMA-3.1-8B Grattafiori et al. (2024). We adopt RLVR-trained models released by SimpleRLZoo Zeng et al. (2025), which train zero-RL models using GRPO on GSM8K and the MATH training set, with correctness reward only, excluding any format-based reward. We compare the pass@$k$ curves of base and zero-RL models on benchmarks of varying difficulty: GSM8K Cobbe et al. (2021), MATH500 Hendrycks et al. (2021), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), AIME24, and AMC23. Additionally, we include the RL model Oat-Zero-7B, based on Qwen-2.5-7B-Base and trained using the Oat-Zero framework, for further comparison Liu et al. (2025b). This model is characterized by strong performance on the challenging AIME24 benchmark, achieving an average of 43.4% points, substantially higher than the base model’s performance of below 10%.
The Effect of RLVR: Increased Likelihood of Correct Samples, Decreased Coverage of Solvable Problems. As shown in Figure 2, we consistently observe a contrasting trend between small and large $k$ values. When $k$ is small (e.g., $k=1$, equivalent to average-case accuracy), RL-trained models outperform their base counterparts. This aligns with the common observation that RL improves performance, suggesting that RLVR makes models significantly more likely to sample correct responses. However, as $k$ increases to the tens or hundreds, base models consistently catch up with RL-trained models across all benchmarks and LLM families without exception. Notably, the pass@$k$ curves of base models exhibit a steeper upward trend compared to their RL-trained versions. Ultimately, the base model surpasses the RL-trained model at sufficiently large $k$, indicating that the coverage of solvable problems by the base model is broader than that of the RL-enhanced version. For example, on the Minerva benchmark with a 32B-sized model, the base model outperforms the RL-trained model by approximately 9% at $k=128$, implying that it can solve around 9% more problems in the validation set.
Figure 3: Oat-Zero for AIME24.
We further examine RL models trained with Oat-Zero. As shown in Figure 3, although the RL model initially demonstrates a strong performance, nearly 30% higher than the base model, it is eventually surpassed by the base model. Based on these results, we conclude that RLVR increases the likelihood of sampling correct responses at low $k$, but narrows the model’s overall coverage. We further analyze the root cause of this phenomenon in Section 4.1.
CoT Case Analysis. We present the sampled correct CoTs from the base model in Figure 18 and Figure 19, manually selected from 2048 samplings for the hardest questions in AIME24. The responses from the base model tend to be long CoTs and exhibit reflective behavior, highlighting the strong reasoning ability inherent in the base model.
Validity of Chain-of-Thought. For mathematical problems, the common evaluation is based solely on the correctness of the final answer, without considering the validity of the intermediate CoT. However, as $k$ increases, the risk of “hacking” may become more pronounced when a model accidentally arrives at the correct answer through an incorrect CoT. To accurately reflect the reasoning ability boundary using pass@$k$, it is important to assess how many of the solved problems result from sampling genuinely correct CoTs, rather than from lucky guesses.
Following Brown et al. (2024), we manually inspect all CoTs that led to correct answers on the most challenging solvable problems in the GSM8k dataset—those with an average accuracy below 5% but above 0%. The base model answered 25 such questions, with 24 containing at least one correct CoT. Similarly, the RL-trained model answered 25 questions, 23 of which included at least one correct CoT. These results suggest that even on the hardest questions in GSM8k, problem-solving primarily results from sampling valid reasoning paths rather than relying on lucky guesses, implying that pass@$k$ at large $k$ values can serve as a relatively accurate indicator of a model’s reasoning boundary.
We also manually check the CoTs for the most challenging AIME24 benchmark. To begin, we introduce a filtering mechanism designed to eliminate easily guessable problems. Specifically, we prompt Qwen-7B-Base to answer questions directly, without using chain-of-thought reasoning, and sample answers multiple times. If a problem can be answered correctly with a low but non-zero probability (e.g., <5%), we consider it guessable and remove it. Problems that can be directly answered correctly with a high probability are retained, as they are likely easier and solvable using valid CoTs. The base and RL model pass@$k$ curves on this filtered AIME24 can be found in Figure 9, showing similar trending with previous results. While this filtering method is heuristic, it proves effective. Applying it to AIME24 (30 questions) results in a subset of 18 questions. We then prompt the models to answer these filtered questions using CoT reasoning. Then we perform a manual inspection of all CoTs that led to correct answers on the hardest problems—those with an average accuracy below 5%. The base model answered 7 such questions, with 5 out of 6 containing at least one correct CoT (excluding one ambiguous case of correctness due to skipped reasoning steps.). Similarly, the RL-trained model answered 6 questions, 4 of which included at least one correct CoT. These results suggest that, even for the hardest questions in the challenging AIME24, base model can sampling valid reasoning paths to solve the problems.
3.2 RLVR for Code Generation
Models and Benchmarks. We adopt the open-source Code-R1 Liu & Zhang (2025) and evaluate its RLVR-trained model, CodeR1-Zero-Qwen2.5-7B, which trains zero-RL models on 12K LeetCode and TACO samples over 832 steps, based on Qwen2.5-7B-Instruct-1M Yang et al. (2025). A binary correctness reward is provided by the compiler based on predefined test cases. For evaluation, Models are assessed on LiveCodeBench v5, comprising 880 problems that span from May 2023 to January 2025 Jain et al. (2024), as well as HumanEval+ and MBPP+ Liu et al. (2023). Since passing all unit tests is nearly impossible to achieve by guesswork, pass@$k$ provides a reliable measure of a model’s reasoning boundary.
Figure 4: RLVR for Coding.
The Effect of RLVR. As shown in Figure 4 and Figure 5 left, the effects of RLVR on three code generation benchmarks exhibit trends that are highly consistent with those observed in mathematical benchmarks. For instance, on LiveCodeBench, the original model achieves a pass@1 score of 23.8%, while the RLVR-trained model achieves 28.1%. However, when sampling 128 times, approximately 50% of coding problems can be solved by the original model, whereas only 42.8% are solved by the RLVR model. Furthermore, we observe that the slope of the original model’s curve remains steep at $k=128$, suggesting that pass@$k$ will continue to increase with larger $k$ values. In contrast, the curve for Code-R1 gradually converges. This suggests that while RLVR shows improvements in single-sample performance, it results in a diminishing coverage boundary and has less potential than the original model.
3.3 RLVR for Visual Reasoning
Models and Benchmarks. In visual reasoning, models must jointly interpret visual and textual inputs to solve complex reasoning problems. This has gained significant attention in the computer vision and multimodal communities since the rise of LLM reasoning Chen et al. (2025); Shen et al. (2025); Zheng et al. (2025). For our experiments, we select mathematical reasoning within visual contexts as a representative task. We use the EasyR1 framework Zheng et al. (2025) to train Qwen-2.5-VL-7B Bai et al. (2025) on Geometry3K Lu et al. (2021), and evaluate its visual reasoning capabilities on filtered MathVista-TestMini Lu et al. (2024) and MathVision-TestMini Wang et al. (2024). Specifically, both MathVista and MathVision contain multiple-choice questions, which can be easily guessed through repeated sampling. Therefore, we remove multiple-choice questions from the evaluation. After filtering, 460 problems remain out of 1000 in MathVista, and 114 problems remain out of 460 in MathVision.
Validity of Chain-of-Thought. To ensure that the increase in coverage is not due to random guessing with incorrect CoT, we manually inspect a subset of the most challenging problems, i.e.those with an average accuracy below 5%. For each inspected problem, we review all CoTs that led to correct answers. We find that for both the original and RL models, 7 out of 8 problems have at least one correct CoT. These results suggest that even on the hardest questions in the filtered set, the increase of coverage primarily results from sampling valid reasoning paths, rather than relying on lucky guesses.
Figure 5: Pass@$k$ curves of base models and zero-RL counterparts. (Left) Code Generation. (Right) Visual Reasoning.
4 Deep Analysis
In this section, we investigate the root cause of the phenomenon where RLVR improves sample efficiency but reduces the reasoning boundary. We also highlight the distinct properties of distillation compared to RLVR. In addition, we design controlled experiments to examine the effects of different RL algorithms.
4.1 Reasoning Patterns Already Present in Base Models
The RLVR Model’s Solvable-Problem Coverage is an Approximate Subset of the Base Model’s Coverage. Experiments in Section 3 surprisingly show that the base model has a larger coverage of solvable problems than the RLVR-trained model. To further investigate, we compare the set of solvable questions for both the base model and its corresponding RL-trained version on AIME24. We find that the set of problems solved by the RL-trained model is nearly a subset of the solvable problems of the base model, as shown in Table 4. A similar trend is observed in coding tasks as shown in Table 5. This raises the natural question: Do all reasoning paths generated by RL-trained models already exist within the output distribution of their base models?
Perplexity Analysis. To answer this question, we utilize the metric perplexity. Given a model $m$, a problem $x$, and a response $\mathbf{Y}=(y_{1},\dots,y_{T})$ (can be generated by the same model, another model, or human), the perplexity is defined as the exponentiated average negative log-likelihood of a sequence:
$\text{PPL}_{m}(\mathbf{Y}|x)=\exp\left(-\frac{1}{T}\sum_{t=1}^{T}\log P(y_{t}|x,y_{1},\dots,y_{t-1})\right),$
We randomly sample two problems from AIME24 and employ Qwen-7B-Base and SimpleRL-Qwen-7B-Base to generate 16 responses for each problem, denoted as $\mathbf{Y}_{\text{base}}$ and $\mathbf{Y}_{\text{RL}}$, respectively. We also let OpenAI-o1 Jaech et al. (2024) generate 8 responses, denoted as $\mathbf{Y}_{\text{GT}}$. As shown in Figure 6 left, the distribution of $\text{PPL}_{\text{Base}}(\mathbf{Y}_{\text{RL}}|x)$ closely matches the lower portion of the $\text{PPL}_{\text{Base}}(\mathbf{Y}_{\text{Base}}|x)$ distribution, corresponding to responses that the base model tends to generate. This suggests that the responses from RL-trained models are highly likely to be generated by the base model. Combining the results from Section 3, we draw the following conclusions:
  • RLVR does not elicit novel reasoning abilities beyond the base model. The trends of pass@$k$ at large values of $k$ and the distribution of perplexity indicate that the reasoning coverage of the RL model remains entirely within that of the base model. All reasoning paths exploited by the RL model already exist within the base model. Thus, RL training does not introduce any fundamentally new reasoning capability, and the RL model remains bounded by the capabilities of its base model.
  • RLVR improves sampling efficiency. Although the reasoning paths in the RL model already exist within the base model, RL training improves pass@1 performance, as shown in Section 3. This indicates that by biasing the output distribution toward high-reward responses, RL increases the likelihood of sampling correct reasoning paths that are already encoded in the base model.
  • RLVR narrows the reasoning boundary. The efficiency gain of RLVR comes at the expense of coverage: pass@$k$ decreases at larger $k$ values compared to the base model. We attribute this to the tendency of RL to reduce output entropy, which limits exploration Yu et al. (2025).
Figure 6: (Left) Perplexity distribution of responses from different sources, evaluated by the base and RL models. The conditioning problem $x$ is omitted in the figure. (Right) Coverage comparison of base, Instruct, RL, and distilled models.
4.2 Distillation Expands the Reasoning Boundary
In addition to direct RL training, another effective approach to improving the reasoning ability of small base models is distillation from a powerful reasoning model such as DeepSeek-R1 Guo et al. (2025). This process is analogous to instruction-following fine-tuning in post-training, converting a base model into an instruct model. However, instead of using short instruction-response pairs, the training data consist of long CoT reasoning traces generated by DeepSeek-R1. Given the limitations of RLVR in expanding reasoning capabilities, it is natural to ask whether distillation exhibits similar behavior. We focus on a representative model, DeepSeek-R1-Distill-Qwen-7B, which distills DeepSeek-R1 into Qwen-2.5-Math-7B. We compare it with the base model Qwen-2.5-Math-7B and its RL-trained counterpart Qwen-2.5-Math-7B-Oat-Zero and include Qwen-2.5-Math-7B-Instruct as an additional baseline. As shown in Figure 6 (right), the pass@$k$ curve of the distilled model is consistently and significantly above that of the base model. This indicates that, unlike RL that is fundamentally bounded by the reasoning capacity of the base model, distillation introduces new reasoning patterns learned from a stronger teacher model. As a result, the distilled model is capable of surpassing the reasoning boundary of the base model.
Figure 7: (Top) Different RL algorithms. (Bottom) Different RL training steps. We use a folded y-axis range to better highlight the details at $k=1$ and $256$. Unfolded version can be found in Figure 8. The detailed values for each point at pass@1 and pass@256 are provided in Table 2 and Table 3.
4.3 Effects of Different RL Algorithms
Experiment Setup. We re-implement popular RL algorithms using the VeRL framework Sheng et al. (2024) for fair comparison, including PPO Schulman et al. (2017), GRPO Shao et al. (2024), Reinforce++ Hu (2025), RLOO Ahmadian et al. (2024), ReMax Li et al. (2023), and DAPO Yu et al. (2025). To assess in-domain and out-of-domain generalization under RLVR, we construct splits from Omni-MATH-Rule, a subset of Omni-MATH Gao et al. (2024) containing verifiable problems. We split it into a training set (2,000 samples) and an in-domain test set (821 samples), and use MATH500 as the out-of-domain benchmark.
As shown in Figure 7 (top), although different RL algorithms exhibit slight variations in both pass@1 and pass@256, these differences are not fundamental. Different RL algorithms yield slightly different $\Delta_{\text{SE}}$ values (i.e., ranging from GRPO’s 43.9 to RLOO’s best 42.6 on the in-domain test set). Furthermore, we observe that $\Delta_{\text{SE}}$ remains consistently above 40 points across different algorithms, highlighting that existing RL methods are still far from achieving optimal sampling efficiency. This suggests that novel RL algorithms or entirely new paradigms may be necessary to approach the upper bound.
We also report several additional observations. First, DAPO achieves slightly higher pass@1 scores across all three datasets; however, its dynamic sampling strategy requires approximately 36×3$\sim$ 6$\times$3 ∼ 6 × more samples per batch during training compared to other algorithms. Moreover, its performance drops significantly at $k=256$. Second, RLOO and Reinforce++ perform consistently well across the entire $k$ range (from 1 to 256), while maintaining efficient training costs, achieving a good balance between effectiveness and efficiency. Third, ReMax shows lower performance at both pass@1 and pass@256. We hypothesize that this is due to its use of the greedy response reward as the advantage baseline, which in the RLVR setting is binary (0 or 1) and highly variable. This likely results in unstable gradient updates during training.
4.4 Asymptotic Effects of RL Training
Based on the setup in Section 4.3, we investigate the effect of the training steps on the asymptotic performance of the model. As shown in Figure 7 (bottom), as RL training progresses, pass@1 on the training set consistently improves from 26.1 to 42.5. However, several observations suggest that longer training may not yield substantial benefits. First, we observe that pass@1 after 450 steps on both in-domain and out-of-domain test sets shows only a marginal improvement compared to 150 steps, which is much smaller than the gain observed in the training set. This suggests that RL training might be overfitting the training set to some degree. Second, we find that as the number of training steps increases, pass@256 decreases across all three sets. Specifically, training for 450 steps results in the lowest pass@256, indicating a minimal reasoning boundary. This is likely because, as RL training progresses, the model’s output entropy and exploration ability decrease.
5 Discussion
@@
In Section 3 and Section 4, we identified key limitations of RLVR in enhancing the LLM reasoning capabilities. In this section, we explore possible underlying factors that may explain why RLVR remains bounded by the reasoning capacity of the base model.
@@
Discussion 2: Priors as a Double-Edged Sword in This Vast Action Space. Since the sampling of responses is guided by the pretrained prior, the policy may struggle to explore new reasoning patterns beyond what the prior already provides. Specifically, in such a complex and highly combinatorial space, most responses generated during training are constrained by the base model’s prior. Any sample deviating from the prior is highly likely to produce invalid or non-sensical outputs, leading to negative reward. As discussed in Section 2.1, policy gradient algorithms aim to maximize the log-likelihood of responses within the prior that receive positive rewards, while minimizing the likelihood of responses outside the prior that receive negative rewards. As a result, the trained policy tends to produce responses already present in the prior, constraining its reasoning ability within the boundaries of the base model. From this perspective, training RL models from a distilled model may temporarily provide a beneficial solution, as distillation helps inject a better prior. In the future, exploration approaches that can explore beyond the confines of the prior in such a vast action space hold promise for creating more powerful reasoning models. Furthermore, alternative paradigms beyond pure RLVR could also be explored to overcome these limitations and enhance the model’s reasoning capabilities.
6 Related Works
Reinforcement Learning for LLM Reasoning. Since the emergence of LLMs, the post-training phase has proven crucial to enhance problem-solving and reasoning abilities Ouyang et al. (2022). This stage typically falls into three main categories: supervised fine-tuning using human-curated or distilled data Wang et al. (2023), self-improvement iteration Zelikman et al. (2022); Gulcehre et al. (2023), and reinforcement learning Ouyang et al. (2022). Previously, a reward model or preferences between responses were employed for reward modeling Ouyang et al. (2022); Rafailov et al. (2023). Recently, Reinforcement Learning with Verifiable Rewards (RLVR) has gained significant traction as a method to improve the reasoning capabilities of LLMs in domains such as mathematics and programming Lambert et al. (2024); Shao et al. (2024). An encouraging landmark work is OpenAI’s o1 model Jaech et al. (2024), which was among the first large-scale applications of RL for reasoning, achieving state-of-the-art results at the time of its release. Following this, Deepseek-R1 Guo et al. (2025) became the first open-weight model to match or surpass the performance of o1. A significant innovation introduced with R1 is the “zero” setting, where reinforcement learning is applied directly to the base LLM, bypassing any intermediate supervised tuning. This approach inspired a wave of open-source efforts to replicate or extend R1’s methodology and improve RL algorithms Zeng et al. (2025); Liu et al. (2025b); Yu et al. (2025); Liu & Zhang (2025). In parallel, reinforcement learning has also gained attention in the multimodal domain, driving advancements in multimodal reasoning Chen et al. (2025); Shen et al. (2025); Zheng et al. (2025).
Although there are many excellent open-source works and algorithmic designs in the field of RLVR, there remains a lack of deep understanding regarding the root effects of RLVR on LLM reasoning abilities and its limitations when starting from the base model. Several studies Liu et al. (2025a); Zhao et al. (2025); Shah et al. (2025) highlight that the reflective behaviors observed in R1-like models actually emerge from the base models, rather than being introduced by RLVR training. Dang et al. (2025) observed a phenomenon similar to our findings: Pass@k deteriorates rapidly and fails to recover with reinforcement learning, but this was seen only in a limited experimental setup with Qwen-2.5-0.5B on GSM8K. More importantly, they did not explore the relationship between the base model and the RL model. In contrast, our paper conducts systematic and rigorous experiments to show that not only reflective behaviors but all reasoning paths are already embedded in the base model. We further demonstrate that RLVR does not elicit any new reasoning abilities beyond the base model.
7 Conclusion
RLVR is widely regarded as a promising approach to enable LLMs to continuously self-improve and acquire novel reasoning capabilities. In this paper, we systematically investigate the reasoning capability boundaries of both base and RLVR models across a wide range of LLM families and benchmarks, using the pass@$k$ metric coupled with manually checked CoTs. Surprisingly, our findings demonstrate that RLVR does not elicit fundamentally new reasoning patterns. Instead, RL primarily enhances the efficiency of LLMs in sampling existing correct reasoning paths encoded in the base model. Consequently, the reasoning boundary remains limited by the base model’s capabilities. Furthermore, our in-depth analysis reveals that current RL algorithms are far from achieving the optimal sampling efficiency, defined by the reasoning boundary of the base model. We also show that distillation plays a significant role in introducing new reasoning patterns and expanding the reasoning boundary. These findings highlight a critical limitation of RLVR in advancing LLM reasoning abilities, suggesting that a new paradigm may be necessary to fully surpass the capabilities of base models.
Appendix A Detailed Experimental Results
The unfolded y-axis version of Figure 7 is provided in Figure 8, and the detailed values of pass@1 and pass@256 are given in Table 2 and Table 3. Indices of solvable problems in AIME24 and LiveCodeBench are provided in Table 4 and Table 5. Pass@$k$ curves in the filtered AIME24 are shown in Figure 9. The effects of different temperature settings are shown in Figure 10.

### v2

@@ A.2 Low-Variance pass@$k$ Estimation
In our experiments, we set $n$ to the largest (i.e., rightmost) $k$ value in the pass@$k$ curves, typically 128, 256, or 1024. For example, in Figure 2, we use $n=128$ for MATH500, Minerva, and GSM8K, and $n=1024$ for AMC23 and AIME24. For the Olympiad benchmark, we set $n=128$ for the Qwen models and $n=1024$ for LLaMA-3.1-8B, due to its relatively lower base model capacity.
Appendix B More Related Works
Reinforcement Learning for LLM Reasoning. Since the emergence of LLMs, the post-training phase has proven crucial to enhance problem solving and reasoning abilities Ouyang et al. (2022). This stage typically falls into three main categories: supervised fine-tuning using human-curated or distilled data Wang et al. (2023), self-improvement iteration Zelikman et al. (2022); Gulcehre et al. (2023), and reinforcement learning Ouyang et al. (2022). Previously, a reward model or preferences between responses were employed for reward modeling Ouyang et al. (2022); Rafailov et al. (2023). Recently, Reinforcement Learning with Verifiable Rewards (RLVR) has gained significant traction as a method to improve the reasoning capabilities of LLMs in domains such as mathematics and programming Lambert et al. (2024); Shao et al. (2024). An encouraging landmark work is OpenAI’s o1 model Jaech et al. (2024), which was among the first large-scale applications of RL for reasoning, achieving state-of-the-art results at the time of its release. Following this, Deepseek-R1 Guo et al. (2025) became the first open-weight model to match or surpass the performance of o1. A significant innovation introduced with R1 is the “zero” setting, where reinforcement learning is applied directly to the base LLM, bypassing any intermediate supervised tuning. This approach inspired a wave of open-source efforts to replicate or extend R1’s methodology and improve RL algorithms Zeng et al. (2025); Liu et al. (2025b); Yu et al. (2025); Liu & Zhang (2025). In parallel, reinforcement learning has also gained attention in the multimodal domain, driving advancements in multimodal reasoning Chen et al. (2025); Shen et al. (2025); Zheng et al. (2025).
Analysis of RLVR. Although there are many excellent open-source works and algorithmic designs in the field of RLVR, there remains a lack of deep understanding regarding the root effects of RLVR on LLM reasoning abilities and its limitations when starting from the base model. Several studies Liu et al. (2025a); Zhao et al. (2025); Shah et al. (2025) highlight that the reflective behaviors observed in R1-like models actually emerge from the base models, rather than being introduced by RLVR training. Dang et al.  Dang et al. (2025) observed a phenomenon similar to our findings: Pass@k deteriorates rapidly and fails to recover with reinforcement learning, but this was seen only in a limited experimental setup with Qwen-2.5-0.5B on GSM8K. More importantly, they did not explore the relationship between the base model and the RL model. In contrast, our paper conducts systematic and rigorous experiments to show that not only reflective behaviors but all reasoning paths are already embedded in the base model. We further demonstrate that RLVR does not elicit any new reasoning abilities beyond the base model.
Appendix C Discussion
@@
In Section 3 and Section 4, we identified key limitations of RLVR in enhancing LLM reasoning capabilities. In this section, we explore possible underlying factors that may explain why RLVR remains bounded by the reasoning capacity of the base model.
@@
Discussion 2: Priors as a Double-Edged Sword in This Vast Action Space. Since the sampling of responses is guided by the pretrained prior, the policy may struggle to explore new reasoning patterns beyond what the prior already provides. Specifically, in such a complex and highly combinatorial space, most responses generated during training are constrained by the base model’s prior. Any sample deviating from the prior is highly likely to produce invalid or non-sensical outputs, leading to negative reward. As discussed in Section 2.1, policy gradient algorithms aim to maximize the log-likelihood of responses within the prior that receive positive rewards, while minimizing the likelihood of responses outside the prior that receive negative rewards. As a result, the trained policy tends to produce responses already present in the prior, constraining its reasoning ability within the boundaries of the base model. From this perspective, training RL models from a distilled model may temporarily provide a beneficial solution, as distillation helps inject a better prior.
Possible Future Work. Developing more efficient exploration strategies may be essential for navigating the vast action space, facilitating the discovery of out-of-prior reasoning patterns and the development of more capable reasoning models. Furthermore, current RLVR frameworks are limited to single-turn interactions with the verifier, lacking the ability to iteratively revise or improve based on environmental feedback. A multi-turn RL agent paradigm—with richer, ongoing interaction with a grounded environment—could enable models to generate novel experiences and learn from them. This agent paradigm has been described as the beginning of an “era of experience” Silver & Sutton (2025).
Appendix D Detailed Experimental Results
D.1 More Results on Mathematics and Coding
Figure 9: More results of SimpleRLZoo on GSM8K and AMC23.
Figure 10: Oat-Zero-7B and DAPO-32B are evaluated on AIME24 and compared against their respective base models.
Figure 11: Coder-R1 on LiveCodeBench.
D.2 Validity of Chain-of-Thought on AIME24
We manually check the CoTs for the most challenging AIME24 benchmark. To begin, we introduce a filtering mechanism designed to eliminate easily guessable problems. Specifically, we prompt Qwen2.5-7B-Base to answer questions directly, without using chain-of-thought reasoning, and sample answers multiple times. If a problem can be answered correctly with a low but non-zero probability (e.g., <5%), we consider it to be guessable and remove it. Problems that can be directly answered correctly with a high probability are retained, as they are likely easier and solvable using valid CoTs.
Figure 12: Pass@$k$ curves of the base and SimpleRLZoo-7B models in the filtered AIME24.
The base and RL model pass@$k$ curves on this filtered AIME24 can be found in Figure 12, showing a similar trending to previous results. Although this filtering method is heuristic, it proves to be effective. Applying it to AIME24 (30 questions) results in a subset of 18 questions. We then prompt the models to answer these filtered questions using CoT reasoning. Then we perform a manual inspection of all CoTs that led to correct answers on the hardest problems – those with an average accuracy below 5%. The base model answered 7 such questions, with 5 out of 6 containing at least one correct CoT (excluding one ambiguous case of correctness due to skipped reasoning steps). Similarly, the RL-trained model answered 6 questions, 4 of which included at least one correct CoT. These results suggest that even for the hardest questions in the challenging AIME24, base model can sample valid reasoning paths to solve the problems.
D.3 Accuracy Distribution Visulization
Figure 13: Accuracy histogram before and after RLVR with SimpleRLZoo models.
D.4 Different RLVR Algorithms
We report several additional observations on different RLVR algorithms in Figure 8. First, DAPO achieves slightly higher pass@1 scores across all three datasets; however, its dynamic sampling strategy requires approximately 36×3$\sim$ 6$\times$3 ∼ 6 × more samples per batch during training compared to other algorithms. Moreover, its performance drops significantly at $k=256$. Second, RLOO and Reinforce++ perform consistently well across the entire $k$ range (from 1 to 256), while maintaining efficient training costs, achieving a good balance between effectiveness and efficiency. Third, ReMax shows lower performance at both pass@1 and pass@256. We hypothesize that this is due to its use of the greedy response reward as the advantage baseline, which in the RLVR setting is binary (0 or 1) and highly variable. This likely results in unstable gradient updates during training.